John is setting up his catering schedule and must determine which events he will be able to cater. Let ei = 1 if event i is selected, 0 if not, for i = 1,...,5. Write the appropriate constraint(s) for each condition. Conditions are independent.

a. Choose no fewer than four projects.
b. If project 1 is chosen, project 3 must be chosen.
c. If project 4 is chosen, project 2 must not be chosen.
d. Projects cost 40, 30, 60, 85, and 28 respectively. The budget is 175.
e. No more than two of projects 2, 3, and 5 can be chosen.
Answer #1

a.

x1+x2+x3+x4+x5 >= 4

b.

x1-x3 <= 0

c.

x4+x2 = 1

d.

40x1+30x2+60x3+85x4+28x5 <= 175

e.

x2+x3+x5 <= 2